Prev Next

SW&IT Meeting Minutes 2019-02-18

Attendees: Andrew, David, Igor, Petr, Stephen, Yang

Outlook

Labor Reporting

Report auto-generated every week, Do not edit this table, use entries in smartsheets.

News from management

Previous Action Items

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

Team reports

MB

  • OC, WPC, management
  • fix dup telemetry files sent to Alta
  • AO s/w effort planning meeting
  • Discuss with AO parties about git migration and benefits, tentatively March
  • obsdev discussions
  • follow up obs workstation upgrade plans, some issue from my VM
  • phone screened 2x new applicants,
next
  • Friday/Monday off
  • github/zenhub
  • obsdev discussions
  • follow up obs workstation upgrade plans
  • SSD 2019 - compile list for Laura (needs your input)

PK

  • GCS redesign - AcqusitionThread, state diagram, GCSState class
  • finally mastered dot (for diagram production), embedded into doxygen documentation
  • lot around locking strategies
next
  • mirror cover meeting (Tuesday 3pm)
  • GCS efforts

DE

  • Rack 3 work
    • Moved VM hosts with no downtime, yay!
    • Moved other less critical machines
    • UPSes to bottom
  • Chasing after Solidworks rep.
  • 4TB drives still available
Next
  • Modify UPS plug with cable on order
  • Networking theory shift
    • Thinking about removing redundant 10Gb switching to use existing switches. This should not really impact reliability.
    • Much cleaner and simpler
    • Will have to math out port numbers; it's possible we'll need more ports in the near-to-medium future, but this will work as a proof of concept.
  • Begin planning Rack 3 population before and during shutdown.

YZ

  • More online training done.
  • AO-CRAVAT:
    • Partially implemented broadcast channel and CSS grid layout, still have trouble with some issues.
    • Continue to investigate solutions to the various problems in web programming.
    • Work on interface polishing more.
  • TCS/IIF:
    • Studying ICE/IIF documents.
    • Need to setup a test environment/cases for the command CorrectModesAO

AC

  • Implementing AGW3 channel into ALH
    • All IT and OAC and most MODS channels read UDF invalid
    • Trouble shooting: Reverted to original tcs.db and ALH-OAC.alhConfig files. Restarted ALH and EPICS-IOC daemon.
    • Corrupt IOC database?
  • Fix and restablish channel access for ALH
    • OAC 5.2 branch to remove quick fix in 5.1
  • Particle Monitor
    • Firmware is severly outdate. Received quote to update it but we will decline service.
next:
  • AGW3 and ALH implmentation
  • build OAC 5.2 on mountain

IS

Last week

  • PIT development, adding LBTI instrument selections
  • Wrap up OT binocular/merging script
  • Met with Michelle a few times to discuss a scheduler for LBTO.

This week

  • Finish PIT development for LBTI
  • Probably start Italian PIT development
  • Continue with new scheduler effort.

SH

  • bought some wireless gear for mountain
  • fedora on to rm507 machine
  • Josh and python
  • nx problems on obs4 and obs2
next:
  • mods and oac alh db.
  • mountain web upgrade fixes
  • tcs-test mount point fixes
  • engineering server
Topic revision: r9 - 18 Feb 2019, IgorSola
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback